OpenStudy (anonymous):

3/4t^-1/4 -1/2t^-3/4 solve for t

OpenStudy (anonymous):

Let's write it down in a better form first , and you'll get ^_^:\[= 3t - t^3/2\] multiply 3t with 2 to get the same denominator : \[=( 6t/2) - t^3/2\] \[= (6t - t^3/2)\] \[= t(6- t^2)/2\] correct me if I'm wrong :)

OpenStudy (anonymous):

i dont think you're wrong

OpenStudy (anonymous):

did you understand the steps? ^_^

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yay ! Good luck n__n

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np ^_^